I need to know how to multiply -2.1 over 2(-0.03) I can't get the correct answer I keep getting a fraction answer on my calculator the answer to this is 35 how did they get that answer??????? help
Answer by richwmiller(17219)

It is 35.
Get rid of the decimals
2*(-.03)=.06
.06*100=6
2.10*100=210
210/6